OpenStudy (anonymous):

Simplify five to the eighth power divided by five to the fourth power using the laws of exponents. Verify your answer by expanding the powers, then dividing. Show your work.

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

\[\frac{ 5^{8} }{ 5^{4} } = \frac{ 5*5*5*5*5*5*5*5 }{ 5*5*5*5 }\] then you can cross out four 5s from the top and bottom \[\frac{ 5*5*5*5*5*5*5*5 }{ 5*5*5*5 } = 5*5*5*5 = 5^{4}\]

OpenStudy (anonymous):

is that it?

jigglypuff314 (jigglypuff314):

well other than I couldn't find the symbols to show to crossing out of the 5s then yeah, that's it :)

OpenStudy (anonymous):

Thanks!